# Is it Really Negative? Evaluating Natural Language Video Localization Performance on Multiple Reliable Videos Pool

### v2

@@ 7.1 Experimental Settings
This section describes the experimental settings of the CroCs and other baselines. More details are shown in Appendix C.
@@
We use standard off-the-shelf video feature extractors without any fine-tuning: VGG Simonyan and Zisserman (2015) feature for Charades-STA; C3D Tran et al. (2015) feature for ActivityNet and TACoS following Wang et al. (2022). We set the filtering hyper-parameter as $t_{tr}=0.9$ for the weakly-supervised contrastive learning. We choose the number of the CroCs hard negative samples to calculate $p^{\prime}(q|v)$ as 100, 200, and 100 and $p^{\prime}(v|q)$ as 50, 100, and 5 for Charades-STA, ActivityNet, and TACoS, respectively. These hyper-parameters are chosen from varying parameter setting experiments as shown in Figure 8. In the hard-negative fine-tuning stage, the learning rates are degraded by multiplying 0.1.
@@
We first select the state-of-the-art models, MMN Wang et al. (2022) and G2L Li et al. (2023), as strong baselines. As the MVMR  task can be formulated as an open-domain question answering (QA) task, we further choose two types of QA models that employ a cross-encoder architecture (i.e., Cross-Encoder Transformer) and a dual-encoder architecture (i.e., Bi-Encoder Transformer), respectively. These architectures are widely adopted in existing NLVL tasks Zhang et al. (2020b); Gao and Xu (2021) due to the closeness of the two tasks.
@@
We evaluate baselines by calculating Rank n@m. It is defined as the percentage of queries having at least one correctly retrieved moment, i.e., IoU $\geq m$, in the top-$n$ derived moments. The existing NLVL studies have reported the results of $m\in\{0.3,0.5,0.7\}$ with $n\in\{1,5\}$ Wang et al. (2022). But, since our MVMR setting searches for massive videos, we report the extended results as $m\in\{0.3,0.5,0.7\}$ with $n\in\{1,5,20,50\}$ for Charades-STA and AcitivityNet and $m\in\{0.3,0.5,0.7\}$ with $n\in\{1,5,20\}$ for TACoS.
7.2 MVMR Results
@@
We evaluate the performance of CroCs and the baseline models on the introduced MVMR datasets. Table 2 shows the experimental results of CroCs and other baselines for our MVMR task. Surprisingly, although CET and BET show quite strong performance on the NLVL task, their performance on the MVMR task notably drops. For examples, CET and MMN show average 34.5 and 29.7 degradation on R@1(IOU0.5) for all MVMR datasets. These outcomes indicate that the current NLVL task performance alone is not enough to verify that a model is well-trained. We reveal that CroCs outperforms all other baselines on MVMR task, and this suggests that our model correctly discriminates features of negative and positive moments. Experiments show that the models on Charades-STA and TACoS significantly improve performance from the weakly-supervised learning since similar queries exist frequently in these datasets and the possibility of false-negatives is high. However, the performance of ActivityNet is attributed to the hard-negative learning since it includes various and complex text queries.
@@
The practical usage scenario for the MVMR should include video retrieval. Therefore, We also conduct MVMR experiments with a video retrieval model, PRMR model Dong et al. (2022), shown as Table 3. We construct a pipeline to filter top-5 logit videos from our MVMR retrieval pool using the PRMR model and solve MVMR task. We use the publicly deployed PRMR model on Charades-STA and ActivityNet datasets. These experiments prove that video retrieval models are helpful to increase MVMR performance, and CroCs still shows superior performance than MMN in this pipeline.
Model R@1 IOU0.3 R@1 IOU0.5 R@5 IOU0.3 R@5 IOU0.5
$\text{MMN}_{\text{Charades-STA}}$ 32.15 27.10 65.97 57.66
$\text{CroCs}_{\text{Charades-STA}}$ 36.64 29.27 68.06 59.78
$\text{MMN}_{\text{ActivityNet}}$ 32.39 25.04 66.84 55.03
$\text{CroCs}_{\text{ActivityNet}}$ 39.25 29.59 71.54 60.02
Table 3: MVMR Experiments with Video Retrieval.
8 Conclusion
@@
In this paper, we propose the Massive Videos Moment Retrieval (MVMR) task, which aims to detect a moment for a natural language query from a massive video set. To stimulate the research, we propose a simple automatic method to construct reliable MVMR datasets and build three MVMR datasets using existing NLVL datasets. We introduce a robust training method called CroCs to distinguish positives from negative distractors effectively in solving MVMR task.
Limitations
Our dataset construction framework automatically defines positive and negative sets using semantic similarity evaluation. Unlike the existing VCMR framework, which overlooks the risk of labeling errors in its automated dataset construction process, our MVMR framework introduces some filtering methods to mitigate labeling errors. However, our method still faces limitations due to its reliance on automatic labeling, which fundamentally carries a risk of labeling inaccuracies. To eliminate the risk of labeling errors thoroughly, a constructed dataset should be reviewed manually through complete enumeration.
Appendix A Datasets Details
Charades-STA
is an extended dataset of action recognition and localization dataset Charades Sigurdsson et al. (2016) by Gao et al. (2017) for video moment retrieval. It is comprised of 5,338 videos and 12,408 query-moment pairs in the training set, and 1,334 videos and 3,720 query-moment pairs in the test set. We report evaluation results on test set in our experiments.
ActivityNet-Captions
is constructed on ActivityNet v1.3 dataset Heilbron et al. (2015), where included videos cover various complex human actions. It is originally designed for video captioning, and recently used into video moment retrieval. It contains 37,417, 17,505, and 17,031 query-moment pairs for training, validation, and testing respectively. We report the evaluation result on val_2 set following the setting of the MMN paper Wang et al. (2022).
TACoS
includes 127 videos selected from the MPII-Cooking dataset. It consists of 18,818 query-moment pairs of various cooking activities in the kitchen annotated by Regneri et al. (2013). It Gao et al. (2017) consists of 10,146, 4,589, and 4,083 query-moment pairs for training, validation and testing, respectively. We report evaluation results on test set in our experiments.
Appendix B Analysis on Filter Methods
@@
We construct reliable MVMR datasets using SimCSE and EMScore. The visualizations of the similarity scores derived by SimCSE and EMScore is shown as Figure 7. For the SimCSE distribution, we select target queries of "person turn a light on", "The person peels a kiwi", and "He is using a push mower to mow the grass" for the Charades-STA, TACoS, and ActivityNet, respectively.
@@
Figure 6 shows the distribution of SimCSE and human-annotated STS16Agirre et al. (2016) dataset scores. STS16 dataset provides text pairs and annotated similarity scores (five-point Likert scale). In the dataset, two texts are regarded as similar when the score is higher than 3. We visualize the distribution of SimCSE and human annotated scores for all text pairs in the dataset, and select $t^{+}_{te}=0.9$ and $t^{-}_{te}=0.5$ as thresholds for the SimCSE filtering to derive MVMR positive and negative candidates, respectively. STS16 dataset includes text pairs that are challenging to distinguish. However, the three datasets utilized in our experiment do not demand similarity measurement at such a complex level. Consequently, we can reliably discern the similarity between queries using the provided thresholds.
@@
Figure 6: The distribution of SimCSE and Human-annotated Scores. The X and Y axes mean the human-annotated and the SimCSE scores, respectively. The width of each graph corresponds to the number of samples.
[…]
Figure 7: The Visulization of the SimCSE and EMScore Similarity Distribution.
Appendix C Implementation Details
@@
We use standard off-the-shelf video feature extractors without any fine-tuning. We use VGG Simonyan and Zisserman (2015) feature for Charades-STA and C3D Tran et al. (2015) feature for ActivityNet and TACoS following the previous work Wang et al. (2022). We train CroCs and MMN on 1 NVIDIA RTX A6000 GPU and early stop by averaging scores of all evaluation metrics.
C.1 CroCs
@@
We choose DistilBERT (Sanh et al., 2019) for the text query encoder following MMN since it is a lightweight model showing significant performance. We calculate the representation of each query $f^{q}\in\mathbb{R}^{d}$ using the global average pooling over all tokens. We adopt the approach of encoding the input video as a 2D temporal moment feature map, inspired by 2D-TAN Zhang et al. (2020b) and MMN. To achieve this, we segment the input video into video clips denoted as $\{c_{i}\}_{i=1}^{l_{c}/k}$, where each clip $c_{i}$ consists of $k$ frames. $l_{c}$ means the total number of frames in the video. The clip-level representations are extracted using a pre-trained visual model (e.g., C3D). By sampling a fixed length with a stride of $\frac{l_{c}}{k\cdot N}$, we obtain $N$ clip-level features. These features are then passed through an FC layer to reduce the dimensionality, resulting in the features $\{f^{v}_{i}\}^{N}_{i=1}$, where $f^{v}_{i}\in\mathbb{R}^{d}$. Utilizing these features, we construct a 2D temporal moment feature map $F\in\mathbb{R}^{N\times N\times d}$ by employing max-pooling as the moment-level feature aggregation method following the baseline (Wang et al., 2022). Additionally, we generate a 2D feature map $F^{\prime}$ of the same size by passing $F$ to 2D Convolution, allowing the representation of moment relations as employed in MMN.
@@
Our convolution network for deriving 2D features is exactly the same as MMN, including the number of sampled clips $N$, the number of 2D conv layers $L$, kernel size, and channels. We set the dimension of the joint feature visual-text space $d=256$, and temperatures $\tau_{m}=\tau_{q}=0.1$. We utilize the pre-trained HuggingFace implementation of DistilBERT Wolf et al. (2019). We set margin $\psi$ as 0.4, 0.3, and 0.1 for Charades-STA, ActivityNet, and TACoS, respectively. We use AdamW Loshchilov and Hutter (2017) optimizer with learning rate of $1\times 10^{-4}$, $8\times 10^{-4}$, $1.5\times 10^{-3}$ for Charades-STA, ActivityNet, and TACoS, respectively. We set $\lambda$ as 0.05 for Charades-STA and TACoS and 0.1 for ActivityNet. We set the reliable true-negatives filtering threshold $t_{tr}=0.9$. We set the number of hard-negative samples to calculate $p^{\prime}(q|v)$ as 100, 200, and 100 and $p^{\prime}(v|q)$ as 50, 100, and 5 for Charades-STA, ActivityNet, and TACoS, respectively. The experimental results for various hyper-parameters for the hard-negative sampling are shown in Figure 8.
@@
Figure 8: Evaluation results for various hyper-parameters. We search various hyper-parameters for the hard negative sampling. We measure R@1 (IoU>m), where m$\in\{0.3,0.5,0.7\}$ to find the best hyper-parameters. In this figure, each $(x,y)$ means the number of negative samples. $x$ and $y$ mean a number of negative samples for $p^{\prime}(v|q)$ (query-to-video) and $p^{\prime}(q|v)$ (query-to-video), respectively. Consequently, we select the number of the CM-ANS samples to calculate $p^{\prime}(q|v)$ as 100, 200, and 100 and $p^{\prime}(v|q)$ as 50, 100, and 5 for Charades-STA, ActivityNet, and TACoS, respectively.
C.2 Baselines
MMN
@@ C.2 Baselines
Wang et al. (2022) adopts a dual-modal encoding design to get video clip and text representations. We utilized the publicly deployed code [^2] https://github.com/MCG-NJU/MMN to implement our own MMN model for our experiment using exactly the same hyper-parameters setting.
G2L
@@
Li et al. (2023) uses a semantically aligned and uniform video grounding framework via geodesic and game theory. We utilized the publicly deployed code [^3] https://github.com/lihxxxxx/G2L?tab=readme-ov-file to implement our own G2L model for our experiment using exactly the same hyper-parameters setting proposed by the paper.
Cross-Encoder Transformer (CET)
@@
is a cross-modal encoding model implemented based on a 6-layered transformer architecture. It is designed to predict the start and end points of a video moment by utilizing a model architecture, which is used in the QA task. We concatenate the pre-extracted video clip features and the text features derived from the Distil-BERT to use as an input of the transformer. To concatenate two extracted features, we should equalize the dimensions of two features with a linear layer as follows:
$\bar{H}^{v}=H^{v}W^{v}$ (7)
$\bar{H}^{q}=H^{q}W^{q}$
where $H^{v}\in\mathbb{R}^{l_{v},d^{v}}$, $H^{q}\in\mathbb{R}^{l_{q},d^{q}}$, $W^{v}\in\mathbb{R}^{d^{v},d}$, $W^{q}\in\mathbb{R}^{d^{q},d}$. $H^{v}$ is the pre-extracted video clip features and $H^{q}$ is the features extracted using Distil-BERT. And then, we derive the representation of each video clip feature from the last layer of the transformer-encoder as follows:
$H^{v,q}=TransformerEnc(\bar{H}^{v}\oplus\bar{H}^{q})$ (8)
@@
$y_{i}^{s}=\left\{\begin{array}[]{ c l }({\frac{x_{e}-i}{x_{e}^{*}-x_{s}^{*}+1}})^{k}&\quad\textrm{if }x_{s}^{*}\leq i\leq x_{e}^{*}\\ 0&\quad\textrm{otherwise}\end{array}\right.\>\>\>\>\>\>$ (10)
$y_{i}^{e}=\left\{\begin{array}[]{ c l }({\frac{i-x_{s}}{x_{e}^{*}-x_{s}^{*}+1}})^{k}&\quad\textrm{if }x_{s}^{*}\leq i\leq x_{e}^{*}\\ 0&\quad\textrm{otherwise}\end{array}\right.\>\>\>\>\>$
We finally use the binary cross-entropy to calculate a loss for the pair of $(p_{i}^{s},y_{i}^{s})$ and $(p_{i}^{e},y_{i}^{e})$ as follows:
$\mathcal{L}^{s}=BCE(p^{s},y^{s})$ (11)
$\mathcal{L}^{e}=BCE(p^{e},y^{e})$
where BCE means the binary cross-entropy. The final loss to train the CET model is defined as follows:
$\mathcal{L}=\frac{\mathcal{L}^{s}+\mathcal{L}^{e}}{2}$ (12)
@@
Bi-Encoder Transformer (BET)
is a dual-modal encoding model that encodes text and video clip features independently. The pre-extracted video clip features are encoded using a 6-layered transformer, and text features are encoded using the Distil-BERT, followed by the average-pooling function and 3-layered MLPs.
$\tilde{H}^{v}=TransformerEnc(\bar{H}^{v})$ (13)
$\tilde{H}^{s}=MLP_{s}(Pool(\bar{H}^{q}))$ (14)
$\tilde{H}^{e}=MLP_{e}(Pool(\bar{H}^{q}))$
@@
$p^{s}=Sigmoid(r*cosine(\tilde{H}^{v},\tilde{H}^{s}))$ (15)
$p^{e}=Sigmoid(r*cosine(\tilde{H}^{v},\tilde{H}^{e}))$
@@
Appendix D Examples of False Negative Videos
Figure 9: Examples of constructed MVMR datasets. We visualize positive and negative video sets for the query “person closes the door" of three constructed datasets. A green solid box means a golden positive moment, and blue solid boxes show moments assigned to videos classified as positive. The underlined queries mean the most similar query described in Section 4.1 (max aggregation).
Illustrative instances of constructed MVMR datasets can be found in Figure 9. This figure shows video instances that have been identified as similar to a specific target query. The target queries correspond to queries of Charades-STA, ActivityNet, and TACoS datasets, sequentially.
Appendix E Human Evaluation for the constructed MVMR Datasets
Figure 10: Human Evaluation Sheet Example.
We conduct a human evaluation on our constructed MVMR datasets to reveal that our method construct MVMR datasets effectively. We recruit crowd workers fluent in English through the university’s online community. The recruited annotators were provided with a detailed description of task definitions, instructions. An example of a sheet we use to conduct a human evaluation is shown in Figure 10.

### v4

@@ 7.1. Experimental Settings
This section describes the experimental settings of the CroCs and other baselines. More details are shown in Appendix B.
@@
We use standard off-the-shelf video feature extractors without any fine-tuning: VGG (Simonyan and Zisserman, 2015) feature for Charades-STA; C3D (Tran et al., 2015) feature for ActivityNet and TACoS following (Wang et al., 2022). We set the filtering hyper-parameter as $t_{tr}=0.9$ for the weakly-supervised potential negative contrastive learning. We choose the number of the CroCs hard negative samples to calculate $p^{\prime}(q|v)$ as 100, 200, and 100 and $p^{\prime}(v|q)$ as 50, 100, and 5 for Charades-STA, ActivityNet, and TACoS, respectively. These hyper-parameters are chosen from varying parameter setting experiments as shown in Figure 9. In the hard-negative fine-tuning stage, the learning rates are degraded by multiplying 0.1.
@@
We first select the state-of-the-art models, MMN (Wang et al., 2022) and G2L (Li et al., 2023), as strong baselines. As the MVMR  task can be formulated as an open-domain question answering (QA) task, we further choose two types of QA models that employ a cross-encoder architecture, Cross-Encoder Transformer (CET), and a dual-encoder architecture, Bi-Encoder Transformer (BET). These architectures are widely adopted in existing VMR tasks due to the closeness of the two tasks.
@@
We evaluate baselines by calculating Rank n@m. It is defined as the percentage of queries having at least one correctly retrieved moment, i.e., IoU $\geq m$, in the top-$n$ derived moments. The existing VMR studies have reported the results of $m\in\{0.3,0.5,0.7\}$ with $n\in\{1,5\}$ (Wang et al., 2022). But, since our MVMR setting searches for massive videos, we report the results of $m\in\{0.5,0.7\}$ with $n\in\{1,5,20,50\}$ for Charades-STA and AcitivityNet and $m\in\{0.5,0.7\}$ with $n\in\{1,5,20\}$ for TACoS.
7.2. MVMR Experimental Results
Overconfidence for the Misinformation.
We evaluate various VMR models (i.g., CET, G2L, and MMN) in both VMR and MVMR settings to determine whether their performance significantly degrades when extending the search coverage to a massive video set. Figure 6 shows that all the baseline models undergo significant performance degradation in the MVMR task. Surprisingly, although CET and G2L show quite strong performance on the VMR task, their performance on the MVMR task notably drops. MMN shows a relatively robust performance than CET and G2L, but it still shows significant performance drops. They tend to be easily distracted by wrong videos and show overconfidence in the misinformation. These outcomes indicate that the current VMR task performance alone is not enough to verify the faithfulness of a VMR model.
Figure 6. VMR vs. MVMR Performance. We report average scores of R1@0.5 and R5@0.5 to reveal the vulnerability of VMR models to misinformation. X and Y axes correspond to the models and the average rank score, respectively.
@@ 7.2. MVMR Experimental Results
Table 2 shows the performance of CroCs and the baseline models on the introduced MVMR datasets. We reveal that CroCs outperforms all other baselines on the MVMR task, and this suggests that our model correctly discriminates features of negative and positive moments. In the Charades-STA and TACoS datasets, the models show significant improvement in performance from the weakly-supervised potential negative learning. These results are attributed to the frequent presence of similar queries in these datasets and the possibility of false negatives. On the other hand, the performance of ActivityNet is attributed to the cross-directional hard-negative learning since it includes various and complex text queries.
Dataset Model R1@0.5 R1@0.7 R5@0.5 R5@0.7 R20@0.5 R20@0.7 R50@0.5 R50@0.7
$\text{MVMR}_{\text{Charades-STA}}$ CET 2.96 2.02 11.79 6.97 32.21 19.91 50.54 33.88
BET 2.99 1.94 10.98 7.00 25.11 16.55 37.86 26.08
G2L 2.78 1.97 10.59 7.36 34.74 21.33 62.72 44.82
MMN 12.78 8.40 33.29 22.36 57.83 41.71 73.09 56.75
CroCs 16.31 10.76 39.05 26.99 62.89 46.39 76.08 60.28
CroCs 17.38 11.44 39.02 27.91 60.98 46.39 76.29 60.76
$\text{MVMR}_{\text{ActivityNet}}$ CET 0.90 0.56 2.32 1.75 4.73 3.87 8.02 6.56
BET 1.16 0.76 2.55 2.03 4.64 3.80 7.10 5.88
G2L 1.32 0.96 6.43 4.55 21.66 14.81 50.07 32.99
MMN 10.71 6.93 25.04 16.55 44.50 30.62 62.09 45.76
CroCs 12.68 7.87 28.47 18.16 46.52 31.12 62.12 44.89
CroCs 15.58 9.51 34.70 22.40 55.58 38.55 70.02 53.20
$\text{MVMR}_{\text{TACoS}}$ CET 5.79 3.89 18.88 12.17 42.53 23.70 - -
BET 5.30 2.97 14.84 8.32 39.76 21.51 - -
G2L 5.82 4.31 21.51 12.29 42.35 22.52 - -
MMN 8.61 5.60 25.84 15.47 43.89 23.80 - -
CroCs 10.75 6.76 27.01 15.28 44.38 25.16 - -
CroCs 10.75 7.30 29.68 16.50 44.77 24.62 - -
Table 2. Experimental Results on MVMR datasets. Bolded and under-lined results indicate the $1^{st}$ and $2^{nd}$ best performance, respectively. CroCs is the model using only weakly-supervised potential negative learning (§6.1). CroCs uses both of two informative sample-weighted learning methods (§6.1 and §6.2). We train all the baselines by three trials and report the averaged MVMR scores.
@@
The practical usage scenario of VMR uses a pipeline integrating video retrieval with video moment retrieval. To analyze in depth the excellence of CroCs, we also conduct MVMR experiments with a pipeline including a video retrieval model, PRMR model (Dong et al., 2022). The experimental results with video retrieval are shown in Table 3. We construct a pipeline to filter top-5 logit videos from our MVMR retrieval pool using the PRMR model and solve the MVMR task. We use the publicly deployed PRMR model on Charades-STA and ActivityNet datasets. These experiments prove that video retrieval models are helpful to increase MVMR performance, and CroCs still shows superior performance than MMN in this pipeline.
Dataset Model R1@0.3 R1@0.5 R5@0.3 R5@0.5
$\text{MVMR}_{\text{Charades-STA}}$ MMN 32.15 27.10 65.97 57.66
CroCs 36.64 29.27 68.06 59.78
$\text{MVMR}_{\text{ActivityNet}}$ MMN 32.39 25.04 66.84 55.03
CroCs 39.25 29.59 71.54 60.02
Table 3. MVMR Experiments with Video Retrieval.
8. Conclusion
@@
In this paper, we propose the Massive Videos Moment Retrieval for Faithfulness Evaluation (MVMR) task, which aims to detect a moment for a natural language query from a massive video set to evaluate the faithfulness of video moment retrieval. To stimulate the research, we propose a simple automatic method to construct trustful MVMR datasets by employing semantic distance verification methods, and finally build three MVMR datasets by extending existing VMR datasets. We further introduce a robust informative sample-weighted learning method called CroCs to distinguish positives from negative distractors effectively in solving the MVMR task.
Acknowledgements.
This work was supported by Adobe Research and LG AI Research. This work was partly supported by Samsung Electronics. This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics & No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University) & No.RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)]. K. Jung is with ASRI, Seoul National University, Korea. The Institute of Engineering Research at Seoul National University provided research facilities for this work.
Appendix A Analysis on Filter Methods
@@
We construct trustful MVMR datasets using SimCSE and EMScore. The visualizations of the results derived by SimCSE and EMScore for the TACoS dataset are shown in Figure 7-(a). We select ”The person peels a kiwi” to show the distribution of SimCSE embeddings.
Figure 7. Qualitative Analysis for Filtering Methods for the TACoS Dataset.
@@
Figure 8 shows the distribution of SimCSE score and human-annotated STS16(Agirre et al., 2016) dataset score. STS16 dataset provides text pairs and annotated similarity scores (five-point Likert scale). In the dataset, two texts are regarded as similar when the score is higher than 3. We visualize the distribution of SimCSE and human annotated scores for all text pairs in the dataset, and select $t^{+}_{te}=0.9$ and $t^{-}_{te}=0.5$ as thresholds for the SimCSE filtering to derive MVMR positive and negative candidates, respectively. STS16 dataset includes text pairs that are challenging to distinguish. However, the three datasets utilized in our experiment do not demand similarity measurement at such a complex level. Consequently, we can reliably discern the similarity between queries using the provided thresholds.
@@
Figure 8. The distribution of SimCSE and Human-annotated Scores. The X and Y axes mean the human-annotated and the SimCSE scores, respectively. The width of each graph corresponds to the number of samples.
Appendix B Implementation Details
@@
We use standard off-the-shelf video feature extractors without any fine-tuning. We use VGG (Simonyan and Zisserman, 2015) feature for Charades-STA and C3D (Tran et al., 2015) feature for ActivityNet and TACoS following the previous work (Wang et al., 2022). We train all VMR models on 1 NVIDIA RTX A6000 GPU and early stop by averaging scores of all evaluation metrics.
B.1. CroCs
@@
We choose DistilBERT (Sanh et al., 2019) for the text query encoder following MMN since it is a lightweight model showing significant performance. We calculate the representation of each query $f^{q}\in\mathbb{R}^{d}$ using the global average pooling over all tokens. We adopt the approach of encoding the input video as a 2D temporal moment feature map, inspired by 2D-TAN (Zhang et al., 2020a) and MMN. To achieve this, we segment the input video into video clips denoted as $\{c_{i}\}_{i=1}^{l_{c}/k}$, where each clip $c_{i}$ consists of $k$ frames. $l_{c}$ means the total number of frames in the video. The clip-level representations are extracted using a pre-trained visual model (e.g., C3D). By sampling a fixed length with a stride of $\frac{l_{c}}{k\cdot N}$, we obtain $N$ clip-level features. These features are then passed through an FC layer to reduce the dimensionality, resulting in the features $\{f^{v}_{i}\}^{N}_{i=1}$, where $f^{v}_{i}\in\mathbb{R}^{d}$. Utilizing these features, we construct a 2D temporal moment feature map $F\in\mathbb{R}^{N\times N\times d}$ by employing max-pooling as the moment-level feature aggregation method following the baseline (Wang et al., 2022). Additionally, we generate a 2D feature map $F^{\prime}$ of the same size by passing $F$ to 2D Convolution, allowing the representation of moment relations as employed in MMN.
@@
Our convolution network for deriving 2D features is exactly the same as MMN, including the number of sampled clips $N$, the number of 2D conv layers $L$, kernel size, and channels. We set the dimension of the joint feature visual-text space $d=256$, and temperatures $\tau_{m}=\tau_{q}=0.1$. We utilize the pre-trained HuggingFace implementation of DistilBERT (Wolf et al., 2019). We set margin $\psi$ as 0.4, 0.3, and 0.1 for Charades-STA, ActivityNet, and TACoS, respectively. We use AdamW (Loshchilov and Hutter, 2017) optimizer with learning rate of $1\times 10^{-4}$, $8\times 10^{-4}$, $1.5\times 10^{-3}$ for Charades-STA, ActivityNet, and TACoS, respectively. We set $\lambda$ as 0.05 for Charades-STA and TACoS and 0.1 for ActivityNet. We set the reliable true-negatives filtering threshold $t_{tr}=0.9$. We set the number of hard-negative samples to calculate $p^{\prime}(q|v)$ as 100, 200, and 100 and $p^{\prime}(v|q)$ as 50, 100, and 5 for Charades-STA, ActivityNet, and TACoS, respectively. The experimental results for various hyper-parameters for the hard-negative sampling are shown in Figure 9.
@@
Figure 9. Evaluation results for various hyper-parameters. We measure Rank 1 (IoU¿m), where m$\in\{0.3,0.5,0.7\}$ to find the best hyper-parameters. Each $(x,y)$ means the number of negative samples. $x$ and $y$ mean a number of negative samples for $p^{\prime}(v|q)$ (query-to-video) and $p^{\prime}(q|v)$ (query-to-video), respectively. We select the number of hard negative samples (§6.2) to calculate $p^{\prime}(q|v)$ as 100, 200, and 100 and $p^{\prime}(v|q)$ as 50, 100, and 5 for Charades-STA, ActivityNet, and TACoS, respectively.
B.2. Baselines
@@
MMN (Wang et al., 2022) adopts a dual-modal encoding design to get video clip and text representations. We utilized the publicly deployed code [^1] https://github.com/MCG-NJU/MMN to implement our own MMN model for our experiment using exactly the same hyper-parameters setting.
@@
G2L (Li et al., 2023) is a sort of stat-of-the-art model, extending MMN architecture. We utilized the publicly deployed code [^2] https://github.com/lihxxxxx/G2L to implement our own G2L model for our experiment using exactly the same hyper-parameters setting.
@@
Cross-Encoder Transformer (CET) is a cross-modal encoding model implemented based on a 6-layered transformer architecture. It is designed to predict the start and end points of a video moment by utilizing a model architecture, which is used in the QA task. We concatenate the pre-extracted video clip features and the text features derived from the Distil-BERT to use as an input of the transformer. To concatenate two extracted features, we should equalize the dimensions of two features with a linear layer as $\bar{H}^{v}=H^{v}W^{v}$ and $\bar{H}^{q}=H^{q}W^{q}$, where $H^{v}\in\mathbb{R}^{l_{v},d^{v}}$, $H^{q}\in\mathbb{R}^{l_{q},d^{q}}$, $W^{v}\in\mathbb{R}^{d^{v},d}$, $W^{q}\in\mathbb{R}^{d^{q},d}$. $H^{v}$ is the pre-extracted video clip features and $H^{q}$ is the features extracted using Distil-BERT. And then, we derive the representation of each video clip feature from the last layer of the transformer-encoder as follows:
(7) $H^{v,q}=TransformerEnc(\bar{H}^{v}\oplus\bar{H}^{q})$
@@
(9) $y_{i}^{s}=\left\{\begin{array}[]{ c l }({\frac{x_{e}-i}{x_{e}^{*}-x_{s}^{*}+1}})^{k}&\quad\textrm{if }x_{s}^{*}\leq i\leq x_{e}^{*}\\ 0&\quad\textrm{otherwise}\end{array}\right.$
$y_{i}^{e}=\left\{\begin{array}[]{ c l }({\frac{i-x_{s}}{x_{e}^{*}-x_{s}^{*}+1}})^{k}&\quad\textrm{if }x_{s}^{*}\leq i\leq x_{e}^{*}\\ 0&\quad\textrm{otherwise}\end{array}\right.$
We finally use the binary cross-entropy to calculate losses for the pair of $(p_{i}^{s},y_{i}^{s})$ and $(p_{i}^{e},y_{i}^{e})$ as $\mathcal{L}^{s}=BCE(p^{s},y^{s})$ and $\mathcal{L}^{e}=BCE(p^{e},y^{e})$, respectively. BCE means the binary cross-entropy. The final loss to train the CET model is defined as follows:
(10) $\mathcal{L}=\frac{\mathcal{L}^{s}+\mathcal{L}^{e}}{2}$
@@
Bi-Encoder Transformer (BET) is a dual-modal encoding model that encodes text and video clip features independently. The pre-extracted video clip features are encoded using a 6-layered transformer, and text features are encoded using the Distil-BERT, followed by the average-pooling function and 3-layered MLPs.
(11) $\tilde{H}^{v}=TransformerEnc(\bar{H}^{v})$
(12) $\tilde{H}^{s}=MLP_{s}(Pool(\bar{H}^{q})),\>\>\>\ \tilde{H}^{e}=MLP_{e}(Pool(\bar{H}^{q}))$
@@
(13) $p^{s}=Sigmoid(r*cosine(\tilde{H}^{v},\tilde{H}^{s})),\>\>\>\ p^{e}=Sigmoid(r*cosine(\tilde{H}^{v},\tilde{H}^{e}))$